Find the values of m and a such that these equations have infinite solution sets.
(i) 4x+my=a
2x+y=4

Answer by nabla(475)

The same line always intersects itself.
Force the 1st equation into similarity with the second:
2x+(m/2)y=a/2
2x+y=4

THUS, we want m/2=1 and a/2=4
So m=2, a=8.